OpenStudy (anonymous):

middle school question: 6+ 3h =-3 um, so I subtracted the six and the three, I subtracted the negative three by 6 and got +3, then when I divide the 3 by 3h I get h=+1, What is it that i'm doing wrong?

OpenStudy (across):

\[6+3h=-3\]\[6-6+3h=-3-6\]\[3h=-9\]\[\frac{3h}{3}=\frac{-9}{3}\]\[h=-3\]

OpenStudy (anonymous):

thank you very much miss!